Assume that a 2 kg of a certain gas with R=0.4305 kJ/kg-K undergoes a process that results in these changes: ΔU=1458.01 kJ, ΔH=2189.13 kJ. What is the temperature change?
1
Expert's answer
2021-07-12T03:36:37-0400

m = 2kg

R = 0.4305 kJ/kg.K

∆U = 1458.01 kJ

∆H = 2189.13 kJ


"∆H = ∆U -mR∆T\\
2189.13 = 1458.01 - 2(0.4305)∆T
\\ 731.12 =0.861∆T\\
∆T = 849.15K"